OpenStudy (anonymous):

Find the hypotenuse of a right triangle with legs of 3 inches and 4 inches.

OpenStudy (whpalmer4):

 Pythagorean theorem says that in a right triangle, the square of the length of the hypotenuse is the sum of the squares of the lengths of the two legs: \[a^2+b^2=c^2\] In your case, \(a = 3\) and \(b = 4\) (inches). Can you find the value of \(c\)? Do not make the common mistake of forgetting to take the square root...